\begin{thebibliography}{99}
%\bibitem{FKY} S.Furuichi, K.Kuriyama and K.Yanagi, Trace inequalities for products of matrices, Linear Alg.Appl., Vol.430(2009),pp.2271-2276.
%\bibitem{AHO} T.Ando, F.Hiai and K.Okubo, Trace inequalities for multiple products of two matrices, Math.Ineq.Appl., Vol.3(2000),pp.307-318.
\bibitem{Furu1} S.Furuichi,A mathematical review of the generalized entropies and their matrix trace inequalities, Proceedings of WEC2007, pp.840-845 (2007).
%\bibitem{JH} C.R. Johnson  and C.J. Hilla, Eigenvalues of words in two positive definite letters, SIAM J. Matrix Anal. Appl.,Vol.23(2002),pp.916-928.
%\bibitem{Bou} J.-C.Bourin, Convexity or concavity inequalities for Hermitian operators, Math. Ineq. Appl. Vol.7(2004),pp.607-620.
\bibitem{MO}     A.W.Marshall and I.Olkin, Inequalities: Theory of majorization and its applications, Academic Press, 1979.
\bibitem{Bha}    R.Bhatia, Matrix Analysis, Springer, 1997.
\bibitem{Gol} S.Golden, Lower bounds for the Helmholtz function, Phys. Rev., Vol.137(1965), pp.B1127-B1128. 
\bibitem{Tho} C.J.Thompson, Inequality with applications in statistical mechanics, J.Math.Phys., Vol.6(1965), pp.1812-1813.

%\bibitem{FYK}    S.Furuichi, K.Yanagi and K.Kuriyama, Fundamental properties of Tsallis relative entropy,                   J.Math.Phys.,Vol.45(2004), pp.4868-4877.
%\bibitem{FYK2} S.Furuichi, K.Yanagi and K.Kuriyama, A note on operator inequalities of Tsallis relative opeartor entropy,  Linear Alg. Appl.,Vol.407(2005), pp.19-31.
%\bibitem{Tsa} C.Tsallis, Possible generalization of Bolzmann-Gibbs statistics, J.Stat.Phys.,Vol.52 (1988), pp.479-487.

%\bibitem{AN} S.Amari and H.Nagaoka, Methods of information geometry, AMS Oxford University Press, 2000.
\bibitem{Furu2} S.Furuichi, Trace inequalities in nonextensive statistical mechanics,  Linear Alg.Appl., Vol.418(2006), pp.821-827.
%\bibitem{HP}    F.Hiai and D.Petz, The Golden-Thompson trace inequality is complemented, Linear Alg. Appl.,Vol.181(1993), pp.153-185.
%\bibitem{HP}    F.Hiai and D.Petz, The Golden-Thompson trace inequality is complemented, Linear Alg. Appl.,Vol.181(1993), pp.153-185.
%\bibitem{Furu3} S.Furuichi, Matrix trace inequalities on Tsallis entropies, J. Inequal. Pure and Appl. Math.,Vol.9(2008),Art.1,7pp.
%\bibitem{Furuta} T.Furuta, Reverse inequalities involving two relative operator entropies and two relative entropies, Linear Alg.Appl., Vol. 403(2005), pp. 24-30.
%\bibitem{Um}  H.Umegaki, Conditional expectation in an operator algebra, IV (entropy and information),Kodai Math.Sem.Rep., Vol.14(1962), pp.59-85.
%\bibitem{BS} V.P.Belavkin and P.Staszewski, C$^*$-algebraic generalization of relative entropy and entropy, Ann. Inst. Henri Poincar\'e, Sec. A, Vol.37(1982), pp.51-58.
%\bibitem{KF} J.I.Fujii and E.Kamei, Relative operator entropy in noncommutative information theory, Math. Japonica, Vol.34(1989), pp.341-348. 
%\bibitem{Sp} W.Specht, Zer Theorie der elementaren Mittel, Math.Z.,Vol.74(1960), pp.91-98.
\end{thebibliography}

\documentclass[11pt]{article}
\usepackage{graphicx}
\usepackage{amsfonts}
\usepackage{theorem}
\newtheorem{Lem}{Lemma}[section]

\newtheorem{The}[Lem]{Theorem}
\newtheorem{Prop}[Lem]{Proposition}
\newtheorem{Cor}[Lem]{Corollary}

\newtheorem{Prob}[Lem]{Problem}

\newtheorem{Con}[Lem]{Conjecture}
\newcommand{\qed}{\hbox{\rule{6pt}{6pt}}}

\begin{document}
\title{A matrix trace inequality and its application}
\author{Shigeru Furuichi$^1$\footnote{E-mail:furuichi@chs.nihon-u.ac.jp} and Minghua Lin$^2$\footnote{E-mail:lin243@uregina.ca}\\
$^1${\small Department of Computer Science and System Analysis,}\\
{\small College of Humanities and Sciences, Nihon University,}\\
{\small 3-25-40, Sakurajyousui, Setagaya-ku, Tokyo, 156-8550, Japan}\\
$^2${\small Department of Mathematics and Statistics,}\\
{\small University of Regina, Regina, Saskatchewan, S4S 0A2, Canada}}
\date{}
\maketitle
{\bf Abstract.} In this short paper, we give a complete and affirmative answer to a conjecture on 
matrix trace inequalities for the sum of positive semidefinite matrices.
We also apply the obtained inequality to derive a kind of generalized Golden-Thompson inequality for
positive semidefinite matrices.
\vspace{3mm}

{\bf Keywords : } Matrix trace inequality, positive semidefinite matrix, majorization and Golden-Thompson inequality

\vspace{3mm}
{\bf 2000 Mathematics Subject Classification : } 15A39 and 15A45
\vspace{3mm}

\section{Introduction}
We give some notations.
The set of all $n \times n$ matrices on the complex field $\mathbb{C}$ is represented by $M(n,\mathbb{C})$.
The set of all $n \times n$ Hermitian matrices is also represented by $M_h(n,\mathbb{C})$.
Moreover the set of all $n \times n$ nonnegative (positive semidefinite) matrices is also represented by $M_+(n,\mathbb{C})$.
Here $X\in M_+(n,\mathbb{C})$ means we have $\langle \phi \vert X \vert \phi \rangle \geq 0$ for any vector
$\vert \phi \rangle \in \mathbb{C}^n$.

The purpose of this short paper is to give the answer to the following conjecture which was given in the paper \cite{Furu1}.
\begin{Con}  {\bf (\cite{Furu1})}\label{con}
For $X,Y\in M_+(n,\mathbb{C})$ and $p\in \mathbb{R}$, the following inequalities hold or not?
\begin{itemize}
\item[(i)] $Tr[(I + X +Y +Y^{1/2}XY^{1/2})^p] \leq Tr[(I+X+Y+XY)^p]$ for $p \geq 1$.
\item[(ii)] $Tr[(I + X +Y +Y^{1/2}XY^{1/2})^p] \geq Tr[(I+X+Y+XY)^p]$ for $0 \leq p \leq 1$.
\end{itemize}
\end{Con}

We firstly note that the matrix $I+X+Y+XY=(I+X)(I+Y)$ is generally not positive semidefinite.
However, the eigenvalues of the matrix $(I+X)(I+Y)$ are same to those of the positive semidefinite matrix
$(I+X)^{1/2}(I+Y)(I+X)^{1/2}$. Therefore the expression $Tr[(I+X+Y+XY)^p]$ always makes sense.

We easily find that the equality for (i) and (ii) in Conjecture \ref{con} holds in the case of  $p=1$. In addition,
the case of $p=2$ was proven by elementary calculations in \cite{Furu1}.

Putting $T=(I+X)^{1/2}$ and $S=Y^{1/2}$, Conjecture \ref{con} can be reformulated by the following problem,
because we have
$Tr[(I+X+Y+XY)^p]=Tr[(T^2+T^2S^2)^p]=Tr[(T^2(I+S^2))^p]=Tr[(T(I+S^2)T)^p]=Tr[(T^2+TS^2T)^p].$

\begin{Prob} \label{prob02}
For $T,S\in M_+(n,\mathbb{C})$ and $p\in \mathbb{R}$, the following inequalities hold or not?
\begin{itemize}
\item[(i)] $ Tr[(T^2+ST^2S)^p] \leq Tr[(T^2+TS^2T)^p]$ for  $p \geq 1$.
\item[(ii)] $ Tr[(T^2+ST^2S)^p] \geq Tr[(T^2+TS^2T)^p]$ for $0 \leq p \leq 1$.
\end{itemize}
\end{Prob}

\section{Main results}
To solve Problem \ref{prob02}, we use the concept of the majorization. See \cite{MO} for the details on the majorization.
Here for $X\in M_h(n,\mathbb{C})$,
$\lambda^{\downarrow }(X)=\left(\lambda_1^{\downarrow}  (X),\cdots,\lambda_n^{\downarrow}  (X) \right)$
 represents the eigenvalues of the Hermitian matrix $X$ in decreasing order,
$\lambda_1^{\downarrow}  (X) \geq \cdots \geq \lambda_n^{\downarrow}  (X)$.
In addition $x \prec y$ means that $x=(x_1,\cdots,x_n)$ is majorized by $y=(y_1,\cdots,y_n)$, if we have
$$\sum_{j=1}^kx_j \leq \sum_{j=1}^k y_j\quad (k=1,\cdots,n-1)$$
and
$$\sum_{j=1}^nx_j =\sum_{j=1}^n y_j.$$

We need the following lemma which can be obtained as a consequence of Ky Fan's maximum principle.
\begin{Lem} {\bf (p.35 in \cite{Bha})}   \label{Lind-lem}
For $A,B \in M_h(n,\mathbb{C})$ and any $k=1,2,\cdots,n$, we have
\begin{equation}
\sum_{j=1}^k \lambda_{j}^{\downarrow}(A+B) \leq \sum_{j=1}^k \lambda_{j}^{\downarrow}(A)+\sum_{j=1}^k \lambda_{j}^{\downarrow}(B).
\end{equation}
\end{Lem}

Then we have the following theorem.
\begin{The} \label{the-01}
For $S,T\in M_+(n,\mathbb{C})$, we have
\begin{equation}   \label{maj_ineq00}
\lambda^{\downarrow }(T^2+ST^2S) \prec \lambda^{\downarrow }(T^2+TS^2T)
\end{equation}
\end{The}

{\it Proof}:
For $S,T\in M_+(n,\mathbb{C})$, we need only to show the following
\begin{equation}   \label{maj_ineq01}
\sum_{j=1}^{k} \lambda_{j}^{\downarrow}  (T^2+ST^2S) \leq \sum_{j=1}^{k} \lambda_{j}^{\downarrow}   (T^2+TS^2T)
\end{equation}
for $k=1,2,\cdots,n-1$,
since we have
$$
\sum_{j=1}^n \lambda_j ^{\downarrow}  (T^2+ST^2S) =\sum_{j=1}^n \lambda_j ^{\downarrow}  (T^2+TS^2T),
$$
which is equivalent to $Tr[T^2+ST^2S]=Tr[T^2+TS^2T]$.

By  Lemma \ref{Lind-lem}, we have
\begin{equation} \label{Lind-ineq01}
2 \sum_{j=1}^k \lambda_{j}^{\downarrow}(X) \leq
\sum_{j=1}^k \lambda_{j}^{\downarrow}\left( X+Y \right)+\sum_{j=1}^k \lambda_{j}^{\downarrow}\left( X-Y \right).
\end{equation}
for $X,Y\in M_h(n,\mathbb{C})$ and any $k=1,2,\cdots,n$.

For $X\in M(n,\mathbb{C})$, the matrices $XX^*$ and $X^*X$ are unitarily similar so that we have $ \lambda_j^{\downarrow}(XX^*) =\lambda_j^{\downarrow}(X^*X) $.
Then we have the following inequality:
\begin{eqnarray*}
2 \sum_{j=1}^{k} \lambda_{j}^{\downarrow}  \left( T^2+TS^2T \right) &=&
\sum_{j=1}^{k} \lambda_{j}^{\downarrow}\left(  T^2+TS^2T \right)  +\sum_{j=1}^{k} \lambda_{j}^{\downarrow}\left( T^2+TS^2T \right)  \\
&=& \sum_{j=1}^{k} \lambda_{j}^{\downarrow}\left( (T+iTS)(T-iST)\right)  +\sum_{j=1}^{k} \lambda_{j}^{\downarrow}\left( (T-iTS)(T+iST)\right)  \\
&=& \sum_{j=1}^{k} \lambda_{j}^{\downarrow}\left(  (T-iST)(T+iTS)\right)  +\sum_{j=1}^{k} \lambda_{j}^{\downarrow}\left(  (T+iST)(T-iTS)\right)  \\
&=& \sum_{j=1}^{k} \lambda_{j}^{\downarrow}\left( T^2+ST^2S+i\left( T^2S-ST^2 \right)\right) +
\sum_{j=1}^{k} \lambda_{j}^{\downarrow}\left(  T^2+ST^2S-i\left( T^2S-ST^2 \right) \right)  \\
&\geq& 2 \sum_{j=1}^{k} \lambda_{j}^{\downarrow} \left( T^2+ST^2S\right),
\end{eqnarray*}
for any $k=1,2,\cdots,n-1$,
by using the inequality (\ref{Lind-ineq01}) for $X=T^2+ST^2S$ and $Y=i(T^2S-ST^2)$.
Thus we have the inequality (\ref{maj_ineq01}) so that the proof is completed.

\hfill \qed

From Theorem \ref{the-01}, we have the following corollary.

\begin{Cor} \label{the01}
For $T,S\in M_+(n,\mathbb{C})$ and $p\in \mathbb{R}$, the following inequalities hold.
\begin{itemize}
\item[(i)] $ Tr[(T^2+ST^2S)^p] \leq Tr[(T^2+TS^2T)^p]$ for  $p \geq 1$.
\item[(ii)] $ Tr[(T^2+ST^2S)^p] \geq Tr[(T^2+TS^2T)^p]$ for $0 \leq p \leq 1$.
\end{itemize}
\end{Cor}

{\it Proof :}
Since $f(x)=x^p,(p\geq 1)$ is convex function and $f(x)=x^p,(0\leq p \leq 1)$ is concave function,
we have the present corollary thanks to Theorem \ref{the-01} and a general property of majorization (See p.40 in \cite{Bha}).
\hfill \qed

As mentioned in Introduction, Corollary \ref{the01} implies the following corollary by putting
$T=(I+X)^{1/2}$ and $S=Y^{1/2}$.

\begin{Cor}  \label{con-cor}
For $X,Y\in M_+(n,\mathbb{C})$ and $p\in \mathbb{R}$, the following inequalities hold.
\begin{itemize}
\item[(i)] $Tr[(I + X +Y +Y^{1/2}XY^{1/2})^p] \leq Tr[(I+X+Y+XY)^p]$ for $p \geq 1$.
\item[(ii)] $Tr[(I + X +Y +Y^{1/2}XY^{1/2})^p] \geq Tr[(I+X+Y+XY)^p]$ for $0 \leq p \leq 1$.
\end{itemize}
\end{Cor}

Thus Conjecture \ref{con} was completely solved with an affirmative answer.


\section{An application}
In this section, we give a kind of one-parameter extension of the famous Golden-Thompson inequality \cite{Gol,Tho} for
positive semidefinite matrices,  applying the obtained result in the previous section.
For this purpose, we denote the generalized exponential function by $\exp_{\nu}(X) \equiv \left( I +\nu X \right) ^{\frac{1}{\nu}}$
for $\nu \in (0,1]$ and $X\in M(n,\mathbb{C})$ such that $Tr[(I +\nu X)^{\frac{1}{\nu}}] \in \mathbb{R}$.
In addition, we use the following inequalities proved in \cite{Furu2}.

\begin{Lem}  {\bf (\cite{Furu2})} \label{GT_ineq_gen_before}
For $X,Y\in M_+(n,\mathbb{C})$, and $\nu \in (0,1]$, we have
\begin{itemize}
\item[(i)]
\begin{equation}   \label{GT_ineq_gen_EQ00}
Tr[  \exp_{\nu}(X+Y)  ] \leq Tr[  \exp_{\nu}(X+Y+\nu Y^{1/2}XY^{1/2})  ].
\end{equation}
\item[(ii)]
\begin{equation}   \label{GT_ineq_gen_EQ}
Tr[\exp_{\nu}(X+Y+\nu XY)] \leq Tr[\exp_{\nu}(X) \exp_{\nu}(Y)].
\end{equation}
\end{itemize}
\end{Lem}

As mentioned in the below of Conjecture \ref{con}, the expression of the left hand side in (\ref{GT_ineq_gen_EQ}) makes also sense, 
since we have $Tr[\exp_{\nu}(X+Y+\nu XY)] = Tr[\left\{ (I+\nu X)^{1/2} (I+\nu Y) (I+\nu X)^{1/2}  \right\}^{\frac{1}{\nu}}] \geq 0$. 

From (i) of Corollary \ref{con-cor} and Lemma \ref{GT_ineq_gen_before}, we have the following proposition.

\begin{Prop}   \label{gen-GT}
For $X,Y\in M_+(n,\mathbb{C})$ and $\nu \in (0,1]$, we have
\begin{equation}\label{gen-GT-ineq}
Tr[  \exp_{\nu}(X+Y)  ] \leq Tr[\exp_{\nu}(X) \exp_{\nu}(Y)].
\end{equation}
\end{Prop}

{\it Proof}:
The right hand side of (\ref{GT_ineq_gen_EQ00}) is bounded from the above by applying (i) of Corollary \ref{con-cor} and
putting $X_1=\nu X$, $Y_1=\nu Y$ and $p=\frac{1}{\nu}$:
\begin{eqnarray*}
Tr\left[\exp_{\nu}(X+Y+\nu Y^{1/2}XY^{1/2})\right] &=& Tr\left[ \left\{ I+\nu (X+Y+\nu Y^{1/2}XY^{1/2}) \right\}^{\frac{1}{\nu}}\right] \\
&=& Tr \left[ (I+X_1+Y_1+Y_1^{1/2}X_1Y_1^{1/2})^p \right]\\
&\leq &  Tr \left[ (I+X_1+Y_1+X_1Y_1)^p \right]\\
&=& Tr \left[ \left\{ I+\nu (X+Y+\nu XY) \right\}^{\frac{1}{\nu}} \right]\\
&=& Tr\left[ \exp_{\nu}(X+Y+\nu XY) \right],
\end{eqnarray*}
which is the left hand side of (\ref{GT_ineq_gen_EQ}). Thus we have the present proposition
thanks to Lemma \ref{GT_ineq_gen_before}.

\hfill \qed

Note that the inequality (\ref{gen-GT-ineq}) can be regarded as a kind of one-parameter extension of the Golden-Thompson inequality for
positive semidefinite matrices $X$ and $Y$.

\section*{Ackowledgements}
We would like to thank the anonymous reviewer for providing valuable comments to improve the manuscript.
The first author (S.F.) was supported in part by the Japanese Ministry of Education, Science, Sports and Culture, Grant-in-Aid for
Encouragement of Young Scientists (B), 20740067

\end{document}